OpenStudy (anonymous):

(6,000 x 10^(-12))(90,000) ----------------------- (30,000)(20 x 10^(-10))

OpenStudy (anonymous):

30,000=3*10^4 20=2*10 combine: 3*10^4*2*10*10^(-10)=3*2*10^(4+1-10)=6*10^(-5) make sense?

OpenStudy (anonymous):

no, whats the answer